OpenStudy (anonymous):

A box of mass 16kg is pulled from rest a distance of 5m across a smooth horizontal floor by a cable inclined at to the horizontal. Find the work done by the tension in the cable given that its magnitude is 25N.

OpenStudy (jamesj):

Work = (Force) x (Displacement in the direction of force) Here, you have the force and the displacement. Hence you just need to calculate.

OpenStudy (jamesj):

Make sense? Talk to me.

OpenStudy (anonymous):

could you give me the answer?? make sense :D

OpenStudy (jamesj):

Well, the force is 25 N and what's the distance?

OpenStudy (anonymous):

5m

OpenStudy (anonymous):

but there is an angle given right?? why don't we include in the calc??

OpenStudy (jamesj):

Ah, what is that angle?

OpenStudy (jamesj):

Are you given a specific value?

OpenStudy (anonymous):

60 degree,and why they give us mass??

OpenStudy (anonymous):

cos 60 is 0.5

OpenStudy (jamesj):

As a distraction. Hence the force in the direction of motion, call it the x-direction, is \[ F_x = F \cos(60) = 25 \frac{1}{2} \] Hence work, W, \[ W = F_x . d = \frac{25}{2} \cdot 5 \ \ \ N \cdot m \]

OpenStudy (jamesj):

Yes, sorry about that.

OpenStudy (anonymous):

thanks :D